Standish Company manufactures consumer products and provided the following information for the month of February:
Units produced 131,400
Standard direct labor hours per unit 0.2 Standard variable overhead rate (per direct labor hour) $3.40
Actual variable overhead costs $88,750
Actual hours worked 26,800
Required:
1. Calculate the variable overhead spending variance using the formula approach. (If you compute the actual variable overhead rate, carry your computations out to five significant digits and round the variance to the nearest dollar.) $______ favorable or unfavorable

2. Calculate the variable overhead efficiency variance using the formula approach. $ ____ favorable or unfavorable

3. What if 26,100 direct labor hours were actually worked in February? What impact would that have had on the variable overhead spending variance?
would be larger
would be smaller
no impact

&

it would be favorable or unfavorable

What impact would that have had on the variable overhead efficiency variance?

would be larger
would be smaller
no impact

&

it would be favorable or unfavorable
Answer #1

information given

Standard hours(SH) = 131400×0.2 = 26280 hours

Standard rate(SR)= $3.40 per hour

Actual hours(AH) = 26800 hours

Actual rate(AR)= 88750/26800 = $3.312 per hour

1.Variable overhead spending variance

= Actual costs − (AH × SR)

= 88750-(26800×3.40)

= $2370 Favorable

2.Variable overhead efficiency variance

= (AH × SR) − (SH × SR)

= (26800×3.40) -(26280×3.4)

= $1768 Unfavorable

information given For 3&4

Standard hours(SH) = 131400×0.2 = 26280 hours

Standard rate(SR)= $3.40 per hour

Actual hours(AH) = 26100 hours

3.Variable overhead spending variance

= Actual costs − (AH × SR)

= 88750-(26100×3.40)

=$ 10 favorable

Impact- overhead spending variance decreased from $2720 favorable to $10 favorable. Variance is smaller.

4.Variable overhead efficiency variance

= (AH × SR) − (SH × SR)

=(26100×3.40) -(26280×3.4)

= $612 unfavorable

Impact- variance decrease from $1768 unfavorable to $612 favorable. Variance is smaller.